Given that the point P(2,-3) lies on the same line as -2x+ky+10=0. find k
Answer by jim_thompson5910(35256)

-2x+ky+10=0

-2(2)+k(-3)+10=0 ... Plug in x = 2 and y = -3

-4 - 3k + 10 = 0

-3k + 6 = 0

-3k = -6

k = -6/(-3)

k = 2